What is the per period rate that corresponds to an EAR of 14.5% if compounding occurs monthly?
Answer #1

EAR = (1-6)*-1 i = stated annual interest rate n = number of compounding periods

i/n is the per period rate, that we need to calculate

I-a(+1) = st!O

+1) = SHIT

1.0113 =(1+=

0.0113 =

Hence, periodic rate = 1.13%